OpenStudy (anonymous):

10 sqrt 96 how do i simplify this?

OpenStudy (anonymous):

Do you mean \[10\sqrt{96}\] or \[\sqrt[10]{96}\] ?

jhonyy9 (jhonyy9):

96=6*16 so sqrt16=4 than will be 4sqrt6 so 10sqrt96=40sqrt6

OpenStudy (anonymous):

factor of 96 = 2x2x2x2x2x3 now sqrt 96 = sqrt 2x2x2x2x2x3 and 10sqrt96 = 10 x 2 x2 sqrt 2x3 = 40 sqrt6 ans